Use the following exchange quotes to help answer problems 18-19. NYSE NASDAQ Bid Ask Bid Ask Company Y mpom $ 127.88 | $ 128.12|$ 127.70 | $ 128.30

Currently you have $120.00 in cash and 1000 shares of Stock Y in your brokerage account. You plan to sell your 1000 shares of Stock Y and use the proceeds to purchase as many shares of Stock X as possible. Assume you cannot borrow or buy partial shares of stock; you will also buy or sell at the best price offered. Based on the quotes provided, the cash balance in your brokerage account after you purchase the shares of Stock X is closest to:

A. $(60.89) B. $59.11 C. $87.95 D. $89.24 E. $106.10
Answer #1

Current Cash = $120

Shares of Company Y will be sold at @$127.88 on NYSE (since higher price is offered) = 1,000*127.88 = $127880

Total Balance = $128,000

Shares of company X will be bought at NASDAQ @150.13 (since sold at lower price)

Number of shares = 128,000/150.13 = 852 shares

Price paid = 852*150.13 = $127,910.76

Cash balance in brokerage account = $89.24

Hence, the answer is D.$89.24